a 3 ft wide sidewalk surrounds a rectangular swimming pool. the perimeter of the pool is 84 ft. the length of the larger rectangle is twice the width of the swimming pool. find the dimensions of the swimming pool
Answer by mananth(16949)

a 3 ft wide sidewalk surrounds a rectangular swimming pool. the perimeter of the pool is 84 ft. the length of the larger rectangle is twice the width of the swimming pool. find the dimensions of the swimming pool
...
perimeter of pool = 84 ft.
width of swimming pool = x
..
length of larger rectangle = 2x
length of the pool = 2x-6 (reduce 3ft from both sides)
...
Perimeter = 2*(l+w)
2*(x+2x-6)=84
2*(3x-6)=84
/2
3x-6=42
+6 +6
3x=48
/3
x= 16 ft. the width of the pool.
length = 2x-6 = 32-6 = 26 feet
POOL width 16 ft length = 26 feet
..
CHECK
16+16+26+26=84
...
m.ananth@hotmail.ca